OpenStudy (anonymous):

can any one tell me that why we use Voltage divider circuit instead of Emitter bias... and also why we use R1 and R2 resistances instead of RB resistance... plzzzzz reply

OpenStudy (kenljw):

With standard Self Bias Network RB = R1||R2 Vbb= Vcc R2/(R1 + R2) = Vcc RB/R1 IB = (Vbb - VBE)/RB = Vcc/R1 - VBE/RB no emitter resistor Your Question is why not just use Vcc and RB for biasing with emitter resistor you can adjust RB << (hFE + 1)RE to eliminate variations in hFE and obtain a stable design IC.

OpenStudy (anonymous):

nd why we use R1 and R2 resistors instead of RB???? i think with the help of RB we might done the biasing then why R1 nd R2..

OpenStudy (kenljw):

The standard variation of hFE is +- 50 % before considering thermal changes in hFE. This vast difference requires the elimination of hFE as a major player is circuit design. Using a RE and adjusting RB, Vbb with R1,R2 is the general method of doing that.

OpenStudy (anonymous):

1)You need only one power supply. 2)Easy to couple/cascade with other stages(check the coupling capacitors C1,C2 and C3 in the attached file.) 3)Simple to get resistors R1 and R2 fabricated in a IC

OpenStudy (kenljw):

Good example of discrete amplifier with couple capacitors to isolate individual amplifier stage biasing. I question 3 above in that its seldom done in IC's. Biasing is usually done with current mirrors, sources, and capacitor coupling is seldom used so as to get a low frequency response.